OpenStudy (anonymous):

Use a matrix to find the solution to the system of equations. -8x-8y=-16 6x-9y=-108

OpenStudy (anonymous):

\[ \begin{bmatrix} -8&-8 \\ 6 & -9 \end{bmatrix} \begin{bmatrix} x \\ y \end{bmatrix} = \begin{bmatrix} -16\\ -108 \end{bmatrix} \]

OpenStudy (anonymous):

so what is the solution and how do i find it?

OpenStudy (abb0t):

1. Divide the first row by -8. 2. multiply the 1st row by 6 3. subtract row 1 from 2nd. 3. divide row 2 by -15 4. subtract R2 from R1

OpenStudy (anonymous):

when you mean row are you talking about the number after the "=" also?

OpenStudy (abb0t):